Dcharger:

Part A: If (62)x = 1, what is the value of x? Explain your answer. Part B: If (60)x = 1, what are the possible values of x? Explain your answer. HELP ASAP

Dcharger:

(6^2 )x = 1

Dcharger:

it should be exponents

Dcharger:

Dcharger:

this one is better

dude:

Ah that helps, thank you

Dcharger:

Yw

Dcharger:

jorellcross press the screenshot its better

Dcharger:

what is annswer for part A and part B

dude:

What you need to remember is that anything to the 0 power will give you 1 Remember the power rules. \(\large (a^x)^y=a^{x\times y}\) So, you want to get the exponents to get to 0 (For the first part) Since there is a zero in the second exponent, any x value will work

Dcharger:

how do you do that

dude:

2 times what will give you 0?

Dcharger:

0

Dcharger:

The rule is that anything multiplied by 0 is equal to 0

dude:

Yeah Thats your answer for A

Dcharger:

but a doesent have 0

Dcharger:

Oh ig that makes sense

Dcharger:

What do I say for part A?

dude:

You can say something along the lines of the possible value of x would be 0 because "The rule is that anything multiplied by 0 is equal to 0" However please do not copy-paste word-for-word what I write

Dcharger:

Yessir i use from my words

Dcharger:

What do I say for part B?

Dcharger:

U workin on it

dude:

Oh I goofed for A Same idea follows, just steps are different \((6^2)^x\) My mistake **Solve \(6^2\) first then find x** This should be \((36)^x\) To get 1, you can now multiply by 0 \(x=0\)

Dcharger:

thats for b or a

dude:

Now for B, since the exponent is already 0 You have, \(6^0=1\) Now you get \((1)^x\) From then you should know what x should be

Dcharger:

x is 0 right

dude:

x=0 for part A, yes

Dcharger:

for part B?

dude:

\(1^x=1\) Do you know how to solve this?

Dcharger:

yes

Dcharger:

1x1

Dcharger:

what are the possible values of x

dude:

Hmm yes that is true but, \(1^1=?\) \(1^2=?\) \(1^3=?\) \(1^4=?\)

Dcharger:

its only 1

dude:

Yeah

dude:

They all work

Dcharger:

Okay so what do i say for possibilities

Dcharger:

11=? 12=? 13=? 14=?

dude:

No, I wrote that for you

Dcharger:

ik

Dcharger:

Since the exponent is already 0 You have, 6^0=1 Now you get 1. The possibilities are 1^1, 1^2, 1^3, 1^4.

Dcharger:

I wrote this

Dcharger:

is this good or bad

dude:

You get 1^x, not just 1

dude:

x can be any whole number, not just 1,2,3,4

Dcharger:

Yo thx for reminding me i was like ill fix that later

Dcharger:

ty